# Krijgslaan, 281-S9, 9000 Gent, Belgium

### v1

@@ Abstract
We analyse a variety of Euclidean saddles in the gravitational path integral, with asymptotic AdS boundary conditions, in a class of Einstein-Scalar-Maxwell models. These include single boundary solutions, usual and wineglass wormholes, as well as more exotic (quasi)-oscillatory saddles. We find several interesting phase transitions between these solutions. The Euclidean wormhole backgrounds can be analytically continued to Lorentzian FLRW universes. Some of them contain an early period of inflation. We delineate the conditions under which they can be the dominant saddles in the gravitational path integral and use them to estimate ratios of probabilities for different cosmological outcomes.
@@ 1 Introduction
In the semiclassical approximation in particular, the gravitational path integral is typically organized as a sum over saddle points of the Euclidean action. In many settings, multiple saddle geometries coexist and compete, with the dominant contribution depending on external parameters such as boundary conditions or the couplings of the specific gravitational theory. Transitions in dominance between distinct saddles give rise to a rich phase structure in the physics of the gravitational systems Hawking:1982dh , analogous but conceptually even richer to thermodynamic phase transitions of non-gravitating local field theory systems. Understanding this competition between saddles is essential for interpreting the gravitational path integral and for determining the different physical regimes of the system at hand.
@@
Among the relevant saddle points that we study in this work are Euclidean wormhole geometries, which connect multiple (in our examples two) asymptotic regions or boundaries and represent intrinsically nonperturbative configurations/solutions of the semiclassical equations of motion (EOMs). In holography and anti-de Sitter/conformal field theory correspondence (AdS/CFT), such saddles lead to the definition of multi-boundary observables and raise sharp questions concerning the factorization of the path integral and the interpretation of bulk gravitational correlations in terms of a dual quantum field theory system Maldacena:2004rf . Their role in the gravitational path integral is therefore closely tied to foundational issues in holography and to the broader problem of defining the gravitational sum over geometries.
@@
A particularly interesting aspect of Euclidean wormhole saddles is their relation, via analytic continuation, to Lorentzian cosmological spacetimes. Different Euclidean saddles may admit continuations leading to qualitatively distinct time-dependent geometries. In particular, while some Euclidean wormhole saddles continue to cosmologies that undergo a big-bang/big-crunch evolution (see Maldacena:2004rf ; Betzios:2019rds ; Betzios:2021fnm ; VanRaamsdonk:2021qgv ; Antonini:2022blk ; Antonini:2022ptt for some early works in the context of holography that emphasized this perspective [^1] Similar wormhole saddles can also be constructed by considering correlated pairs of operators in two CFT copies Betzios:2023obs ; Maloney:2025tnn ; VanRaamsdonk:2026tnv or by appropriate pinching limits of higher genus Riemann surfaces in $3d$ gravity Belin:2025ako .), there exist others that can remarkably give rise to inflationary cosmological spacetimes Betzios:2024oli ; Betzios:2024zhf  [^2] We should emphasize here that this construction is different from quantum tunnelling/Coleman de-Lucia processes DeAlwis:2019rxg or the “bag-of-gold” spacetimes behind black hole horizons Fu:2019oyc , where a portion of Lorentzian de Sitter could appear inside Lorentzian anti-de Sitter. These later cases seem to suffer from a version of the information paradox Freivogel:2005qh , whereby a pure state has to evolve into a mixed state.. The competition between Euclidean saddles in the path integral thus translates, after continuation, into a competition between distinct cosmological outcomes. Clarifying how this semiclassical phase structure manifests itself in Lorentzian signature provides a new perspective on the emergence of cosmological spacetimes from holography.
@@
In this work, in section 2, we start with a general discussion on the relation between different classes of Euclidean saddle points of the gravitational path integral and how they can be used in a cosmological setting to define normalised probabilities of different outcomes of the analytically continued cosmologies at the semi-classical level. Due to various issues inherent to the no-boundary (NB) proposal of Hartle and Hawking Hartle:1983ai ; Lehners:2023yrj ; Maldacena:2024uhs , we are driven to consider the possibility that well defined non-trivial cosmological states/wavefunctions and probabilities can only exist in models that admit asymptotically AdS gravitational saddles Betzios:2024oli ; Betzios:2024zhf , whereby different dominant saddles correspond to distinct cosmological outcomes, including crunching and expanding universes.
@@ 2 Cosmological wavefunctions, transition amplitudes and probabilities
In this section, we discuss the cosmological interpretation of our results [^4] More details that include perturbations of fields will be presented in future work Betzios2025toappear ., noticing that Euclidean wormhole backgrounds can be analytically continued to Lorentzian FLRW cosmologies Maldacena:2004rf ; Betzios:2017krj ; Betzios:2019rds ; Betzios:2024oli ; Betzios:2024zhf ; VanRaamsdonk:2021qgv ; Antonini:2022blk ; Antonini:2022ptt .
We first start with the Hartle-Hawking wavefunction Hartle:1983ai , that can be computed using the Euclidean path integral
@@
This path integral is over compact Euclidean geometries in the Euclidean past (“no-boundary”), so that there is only a single boundary on which we define a bulk gravity state/wavefunction (here we take it to be at $\tau=0$—at the “birth” of the Lorentzian Universe). We can then evolve this wavefunction to a real Lorentzian time $t$. The resulting complete gravitational path integral is described at the semi-classical level by a complex geometry that interpolates between Euclidean and Lorentzian de-Sitter (assuming $V(\phi)>0$). While being a quite elegant and beautiful construction, the no-boundary proposal Hartle:1983ai is plagued by notorious issues, reviewed in Quantum:Cosmology ; Lehners:2023yrj ; Maldacena:2024uhs ; Betzios:2024oli , the most important being that it is at clash with inflationary theory, predicting the smallest possible number of inflationary e-folds. Additionally, it cannot be derived or justified within any known consistent model of quantum gravity such as holography or string theory. In a recent work Abdalla:2026mxn , even the original interpretation of the gravitational path integral in eq. (1) as a wavefunction by Hartle and Hawking came at question, and upon using a standard normalization of gravitational transition amplitudes (i.e. Born’s rule)
@@
the authors were led to the remarkable conclusion that classically distinct data $|h_{ij},\phi\rangle$, actually define nearly parallel states of the gravitational Hilbert space up to non-perturbatively small differences in the Newton’s constant $G_{N}$. This would imply that the (normalised) transition probability from one closed universe state to any other is also approximately one, which seems to disagree with even the most basic quantum mechanical rules and experiments. Given this lack of predictability, the authors of  Abdalla:2026mxn also considered the possibility of projecting out the no-boundary state (for example by redefining the rules of the gravitational path-integral). One is thus naturally led to consider alternative proposals for the wavefunction of the universe that incorporate the presence of asymptotic boundaries—one such possibility being that EAdS wormholes such as those in Maldacena:2004rf ; Betzios:2019rds ; Betzios:2021fnm ; VanRaamsdonk:2021qgv ; Antonini:2022blk ; Antonini:2022ptt ; Betzios:2024oli ; Betzios:2024zhf , are the semiclassical saddles that should be used to define non-trivial bulk gravitational states, see Betzios:2024oli for more details on this proposal in the context of inflationary cosmology.
@@
At this point we should emphasize that due to different asymptotic conditions in the Euclidean past, one cannot really compare the no-boundary with the wormhole saddles in the gravitational path integral—they actually seem to descend from different theories of quantum gravity [^6] Unless there is a way to make sense of the sum over the class of all boundary conditions including the cases of not even having a boundary as well as having sources or no sources. This would correspond to a sum over “states” of some sort of “metatheory” and transitions between such “states” do not seem possible to happen for any physical system that we know so far.. An alternative way to phrase this is that the space of all asymptotic possibilities/configurations can be split into a “superselection sector sum” $\mathcal{H}=\cup_{\rm asym.}\mathcal{H}_{\rm asym.}$, where with “asym.” we label classes with inherently different asymptotics i.e AdS vs. flat boundaries vs. no-boundaries and different sources, for which no finite energy operation can induce cross-transitions between them. This leads to an equivalent conclusion: different asymptotics define different theories of quantum gravity when there is no physical operation that can induce transitions between them (see Banks:2025nfe ; Sen:2025bmj for similar ideas). From those possibilities, and within our current knowledge of microscopic models of quantum gravity, the asymptotic AdS wormhole proposal is the only one admitting a well defined UV completion with a well defined underlying Hilbert space of states encoded in the Holographic dual QFT.
This also means that one needs to modify slightly and enrich the notion of probabilities in this setting compared to the recent analyses of the no-boundary proposal, for which the Hilbert space of closed Universes is found to be one-dimensional. As emphasized in Betzios:2022oef one has on the one hand the (UV) asymptotic AdS boundary ($AB$) Hilbert space of states $\mathcal{H}_{AB}$ (descending from the Hilbert space of the dual QFT and leading to different vacuum expectation values (VEVs) for bulk fields in a region where the geometry and fields become non fluctuating) and on the other hand the interior (bulk) boundary Hilbert space of states $\mathcal{H}_{IB}$ (IBIB) (which is defined upon slicing a geometry on a locus where the metric and fields are dynamical and fluctuate and where the gravitational constraints are most important). Any mapping between these two Hilbert spaces of states is bound to be highly non-trivial and rests upon the precise definition and operational details of holography. The meaningful probabilities that one can define in this setting and which are also relevant for cosmology, are computed using transition amplitudes from an $AB$-state to some $IB$-state
@@
In practice, they can be computed at the semiclassical level using ratios of the on-shell actions of connected (half) wormhole geometries with wavefunctions at $\tau=0$ corresponding to the various $IB_{i}$ states. More generally, if one wishes to compute probabilities that involve later real times $t$, one has to consider complex geometries, part of which corresponds to the (half)-wormhole and a part of which depends on the specific inflationary model/region of the scalar potential (we leave this more complete analysis for the future Betzios2025toappear ).
In the model we analyze in this work, we find that for certain choices of UV BCs. ($AB$) a wineglass (half) wormhole manifold dominates, while for others a simple connected (half) wormhole dominates. This simply means that in the first case, it is much more probable that the Universe gets nucleated at $\tau=0$ and then starts to inflate, while in the second case it is more probable that the Universe will undergo a big-Crunch collapse. From the $AB$-cases for which inflation is the most probable outcome, the dominant initial condition ($IB_{i}$) for the state of the Universe is determined from the classical solution to the bulk EOMs [^8] As mentioned previously there are always off-shell or complex contributions for transitions to various $IB_{i}$ states, but these are all suppressed at the semi-classical level, due to the existence of a real classical saddle. - in the wineglass case, this is found to be high up in the inflationary plateau of the effective potential $V(\phi)$, giving further credence to the proposal of Betzios:2024oli ; Betzios:2024zhf as providing a good initial state for the beginning of an inflationary Universe. The basic question remaining is to determine the conditions under which “wineglass” wormholes (or quasi-oscillatory wormholes of the type that we describe in section 4) can be the dominant connected saddles of the gravitational path integral, rendering inflation as the most probable outcome for the early Universe cosmology.
@@ 3 The Einstein-scalar-radiation model
Our starting point is the Einstein–Hilbert action with a minimally coupled scalar field $\phi$ (the inflaton), in the presence of gauge/radiation fields which play an important role in shaping the solutions of interest. This is a quite general model that forms a subsector of the Standard Model in the case that the scalar inflaton is identified with the Higgs boson Bezrukov:2007ep ; Barvinsky:2008ia ; DeSimone:2008ei ; Barbon:2009ya ; Barvinsky:2009fy ; Barvinsky:2009ii ; Bezrukov:2010jz ; Bezrukov:2014ipa ; Betzios:2024zhf ; Gialamas:2025kef . We shall work in four dimensions motivated from phenomenological purposes, but similar solutions can be found in an arbitrary number of dimensions.
@@
where $\mathcal{R}$ denotes the scalar curvature, $V(\phi)$ is the scalar field potential, $\kappa=1/M_{\rm Pl}^{2}$ and $\mathcal{S}_{\rm GHY}$ is the Gibbons-Hawking-York boundary term which has been added to render the variational problem well defined. Higher–order curvature corrections and possible non-minimal couplings between gravity and the scalar field are also expected to be generated at the quantum level. Such terms could, in principle, have a significant impact on the inflationary dynamics of the system, for example by flattening the scalar potential and modifying the inflationary trajectory leading to different inflationary parameters (see Bezrukov:2010jz ; Bezrukov:2014ipa ; Gialamas:2025kef and Betzios:2024zhf for an application in the present context). Nevertheless, in what follows we shall keep the discussion and analysis general within our minimal setup, since it can be applied to any particular inflationary model.
@@
The dominance of $SO(4)$-symmetric Euclidean configurations in the path integral Coleman:1985rnk , which reflects the universe’s observed isotropy and homogeneity, motivates the use of the closed (spherical) case, $k=1$ , which we shall adopt in the rest of this work. An important aspect to consider is the sign of the electromagnetic energy density $\tilde{\rho}^{E}_{\rm rad}$. To obtain the Euclidean wormhole solutions for the cosmological purposes of interest Betzios:2024oli , it is necessary that $\tilde{\rho}^{E}_{\rm rad}<0\,$. This requirement becomes clear from the third equation in (9), where the curvature term $1/a^{2}$ tends to reduce the scale factor, whereas the term $\tilde{\rho}^{E}_{\rm rad}/a^{4}$, with $\tilde{\rho}^{E}_{\rm rad}<0$ , contributes to its expansion. The competition between these terms gives rise to the wormhole neck and supports the subsequent accelerated expansion of the universe. The condition $\tilde{\rho}^{E}_{\rm rad}<0\,$ can be naturally realized if the magnetic field contribution dominates over the electric one in the early universe Betzios:2024zhf  [^9] A similar mechanism can also work in braneworld models such as the ones studied in Antonini:2024bbm and it would be interesting to extend our analysis in this direction.. In particular, the Euclidean Lagrangian of the electromagnetic field in the $U(1)$ case is $\mathcal{L}_{\rm rad}^{E}=F_{\mu\nu}^{E}F^{E\,\mu\nu}/4=(E^{2}+B^{2})/2\,$, which, through the corresponding electromagnetic energy-momentum tensor, ${T^{E}_{\rm rad}}^{\mu}_{\,\,\nu}$ , leads to a Euclidean radiation energy density of the form
@@
A similar mechanism can be demonstrated for axions (or two form) fields instead of gauge (or one form) fields, see for example Betzios:2024oli . This again stems from the fact that the axionic contribution to the energy density $\rho_{\rm axion}$ arising from a three form flux $H_{ijk}$ is negative in the Euclidean signature, since it is always of the magnetic type (no $\tau$ component). A final type of mechanism giving rise to Euclidean wormhole saddles is to consider slices that have negative curvature i.e. k=1k=-1\, (combined now with dominant electric radiation density). The slices should be then compactified (in analogy to higher genus Riemann surfaces in two dimensions), if the wormhole interpretation of the backgrounds is to be retained Betzios:2019rds .

### v2

@@ Abstract
We analyse a variety of Euclidean saddles in the gravitational path integral, with asymptotic AdS boundary conditions, in a class of Einstein-Scalar-Maxwell models. These include single boundary solutions, usual and wineglass wormholes, as well as more exotic (quasi)-oscillatory saddles. Our construction shows how an unbound number of oscillations gets tamed, when flat directions of the potential get lifted. We find several interesting phase transitions between these solutions. The Euclidean wormhole backgrounds can be analytically continued to Lorentzian FLRW universes. Some of them contain an early period of inflation. We delineate the conditions under which they can be the dominant saddles in the gravitational path integral and use them to estimate ratios of probabilities for different cosmological outcomes.
@@ 1 Introduction
In the semiclassical approximation in particular, the gravitational path integral is typically organized as a sum over saddle points of the Euclidean action. In many settings, multiple saddle geometries coexist and compete, with the dominant contribution depending on external parameters such as boundary conditions or the couplings of the specific gravitational theory. Transitions in dominance between distinct saddles give rise to a rich phase structure in the physics of the gravitational systems Hawking and Page (1983), analogous but conceptually even richer to thermodynamic phase transitions of non-gravitating local field theory systems. Understanding this competition between saddles is essential for interpreting the gravitational path integral and for determining the different physical regimes of the system at hand.
@@
Among the relevant saddle points that we study in this work are Euclidean wormhole geometries, which connect multiple (in our examples two) asymptotic regions or boundaries and represent intrinsically nonperturbative configurations/solutions of the semiclassical equations of motion (EOMs). In holography and anti-de Sitter/conformal field theory correspondence (AdS/CFT), such saddles lead to the definition of multi-boundary observables and raise sharp questions concerning the factorization of the path integral and the interpretation of bulk gravitational correlations in terms of a dual quantum field theory system Maldacena and Maoz (2004). Their role in the gravitational path integral is therefore closely tied to foundational issues in holography and to the broader problem of defining the gravitational sum over geometries.
@@
A particularly interesting aspect of Euclidean wormhole saddles is their relation, via analytic continuation, to Lorentzian cosmological spacetimes. Different Euclidean saddles may admit continuations leading to qualitatively distinct time-dependent geometries. In particular, while some Euclidean wormhole saddles continue to cosmologies that undergo a big-bang/big-crunch evolution, see Maldacena and Maoz (2004); Betzios et al. (2019, 2022b); Van Raamsdonk (2022); Antonini et al. (2024, 2023) for some early works in the context of holography that emphasized this perspective [^1] Similar wormhole saddles can also be constructed by considering correlated pairs of operators in two CFT copies Betzios and Papadoulaki (2024b); Maloney et al. (2025); Van Raamsdonk and Vilar López (2026) or by appropriate pinching limits of higher genus Riemann surfaces in $3d$ gravity Belin and de Boer (2025)., there exist others that can remarkably give rise to inflationary cosmological spacetimes Betzios and Papadoulaki (2024a); Betzios et al. (2025) [^2] We should emphasize here that this construction is different from quantum tunnelling/Coleman de-Lucia processes De Alwis et al. (2020) or the “bag-of-gold” spacetimes behind black hole horizons Fu and Marolf (2019), where a portion of Lorentzian de Sitter could appear inside Lorentzian anti-de Sitter. These later cases seem to suffer from a version of the information paradox Freivogel et al. (2006), whereby a pure state has to evolve into a mixed state. - see also Lavrelashvili et al. (1988) for an early incarnation of this idea in asymptotically flat space. The competition between Euclidean saddles in the path integral thus translates, after continuation, into a competition between distinct cosmological outcomes. Clarifying how this semiclassical phase structure manifests itself in Lorentzian signature provides a new perspective on the emergence of cosmological spacetimes from holography.
@@
In this work, in section 2, we start with a general discussion on the relation between different classes of Euclidean saddle points of the gravitational path integral and how they can be used in a cosmological setting to define normalised probabilities of different outcomes of the analytically continued cosmologies at the semi-classical level. Due to various issues inherent to the no-boundary (NB) proposal of Hartle and Hawking Hartle and Hawking (1983); Lehners (2023); Maldacena (2024), we are driven to consider the possibility that well defined non-trivial cosmological states/wavefunctions and probabilities can only exist in models that admit asymptotically AdS gravitational saddles Betzios and Papadoulaki (2024a); Betzios et al. (2025), whereby different dominant saddles correspond to distinct cosmological outcomes, including crunching and expanding universes.
@@ 2 Cosmological wavefunctions, transition amplitudes and probabilities
In this section, we discuss the cosmological interpretation of our results [^4] More details that include perturbations of fields will be presented in future work Betzios et al. (To appear)., noticing that Euclidean wormhole backgrounds can be analytically continued to Lorentzian FLRW cosmologies Maldacena and Maoz (2004); Betzios et al. (2018, 2019); Betzios and Papadoulaki (2024a); Betzios et al. (2025); Van Raamsdonk (2022); Antonini et al. (2024, 2023).
We first start with the Hartle-Hawking wavefunction Hartle and Hawking (1983), that can be computed using the Euclidean path integral
@@
This path integral is over compact Euclidean geometries in the Euclidean past (“no-boundary”), so that there is only a single boundary on which we define a bulk gravity state/wavefunction (here we take it to be at $\tau=0$—at the “birth” of the Lorentzian Universe). We can then evolve this wavefunction to a real Lorentzian time $t$. The resulting complete gravitational path integral is described at the semi-classical level by a complex geometry that interpolates between Euclidean and Lorentzian de-Sitter (assuming $V(\phi)>0$). While being a quite elegant and beautiful construction, the no-boundary proposal Hartle and Hawking (1983) is plagued by notorious issues, reviewed in Coleman et al. (1991); Lehners (2023); Maldacena (2024); Betzios and Papadoulaki (2024a), the most important being that it is at clash with inflationary theory, predicting the smallest possible number of inflationary e-folds. Additionally, it cannot be derived or justified within any known consistent model of quantum gravity such as holography or string theory. In a recent work Abdalla et al. (2026), even the original interpretation of the gravitational path integral in eq. (1) as a wavefunction by Hartle and Hawking came at question, and upon using a standard normalization of gravitational transition amplitudes (i.e. Born’s rule)
@@
the authors were led to the remarkable conclusion that classically distinct data $|h_{ij},\phi\rangle$, actually define nearly parallel states of the gravitational Hilbert space up to non-perturbatively small differences in the Newton’s constant $G_{N}$. This would imply that the (normalised) transition probability from one closed universe state to any other is also approximately one, which seems to disagree with even the most basic quantum mechanical rules and experiments. Given this lack of predictability, the authors of  Abdalla et al. (2026) also considered the possibility of projecting out the no-boundary state (for example by redefining the rules of the gravitational path-integral). One is thus naturally led to consider alternative proposals for the wavefunction of the universe that incorporate the presence of asymptotic boundaries—one such possibility being that EAdS wormholes such as those in Maldacena and Maoz (2004); Betzios et al. (2019, 2022b); Van Raamsdonk (2022); Antonini et al. (2024, 2023); Betzios and Papadoulaki (2024a); Betzios et al. (2025), are the semiclassical saddles that should be used to define non-trivial bulk gravitational states, see Betzios and Papadoulaki (2024a) for more details on this proposal in the context of inflationary cosmology.
@@
At this point we should emphasize that due to different asymptotic conditions in the Euclidean past, one cannot really compare the no-boundary with the wormhole saddles in the gravitational path integral—they actually seem to descend from different theories of quantum gravity [^6] Unless there is a way to make sense of the sum over the class of all boundary conditions including the cases of not even having a boundary as well as having sources or no sources. This would correspond to a sum over “states” of some sort of “metatheory” and transitions between such “states” do not seem possible to happen for any physical system that we know so far.. An alternative way to phrase this is that the space of all asymptotic possibilities/configurations can be split into a “superselection sector sum” $\mathcal{H}=\cup_{\rm asym.}\mathcal{H}_{\rm asym.}$, where with “asym.” we label classes with inherently different asymptotics i.e AdS vs. flat boundaries vs. no-boundaries and different sources, for which no finite energy operation can induce cross-transitions between them. This leads to an equivalent conclusion: different asymptotics define different theories of quantum gravity when there is no physical operation that can induce transitions between them (see Banks (2025); Sen (2025) for similar ideas). From those possibilities, and within our current knowledge of microscopic models of quantum gravity, the asymptotic AdS wormhole proposal is the only one admitting a well defined UV completion with a well defined underlying Hilbert space of states encoded in the Holographic dual QFT.
This also means that one needs to modify slightly and enrich the notion of probabilities in this setting compared to the recent analyses of the no-boundary proposal, for which the Hilbert space of closed Universes is found to be one-dimensional. As emphasized in Betzios et al. (2022a) one has on the one hand the (UV) asymptotic AdS boundary ($AB$) Hilbert space of states $\mathcal{H}_{AB}$ (descending from the Hilbert space of the dual QFT and leading to different vacuum expectation values (VEVs) for bulk fields in a region where the geometry and fields become non fluctuating) and on the other hand the interior (bulk) boundary Hilbert space of states $\mathcal{H}_{IB}$ (IBIB) (which is defined upon slicing a geometry on a locus where the metric and fields are dynamical and fluctuate and where the gravitational constraints are most important). Any mapping between these two Hilbert spaces of states is bound to be highly non-trivial and rests upon the precise definition and operational details of holography. The meaningful probabilities that one can define in this setting and which are also relevant for cosmology, are computed using transition amplitudes from an $AB$-state to some $IB$-state
@@
In practice, they can be computed at the semiclassical level using ratios of the on-shell actions of connected (half) wormhole geometries with wavefunctions at $\tau=0$ corresponding to the various $IB_{i}$ states. More generally, if one wishes to compute probabilities that involve later real times $t$, one has to consider complex geometries, part of which corresponds to the (half)-wormhole and a part of which depends on the specific inflationary model/region of the scalar potential (we leave this more complete analysis for the future Betzios et al. (To appear)).
In the model we analyze in this work, we find that for certain choices of UV BCs. ($AB$) a wineglass (half) wormhole manifold dominates, while for others a simple connected (half) wormhole dominates. This simply means that in the first case, it is much more probable that the Universe gets nucleated at $\tau=0$ and then starts to inflate, while in the second case it is more probable that the Universe will undergo a big-Crunch collapse. From the $AB$-cases for which inflation is the most probable outcome, the dominant initial condition ($IB_{i}$) for the state of the Universe is determined from the classical solution to the bulk EOMs [^8] As mentioned previously there are always off-shell or complex contributions for transitions to various $IB_{i}$ states, but these are all suppressed at the semi-classical level, due to the existence of a real classical saddle. - in the wineglass case, this is found to be high up in the inflationary plateau of the effective potential $V(\phi)$, giving further credence to the proposal of Betzios and Papadoulaki (2024a); Betzios et al. (2025) as providing a good initial state for the beginning of an inflationary Universe. The basic question remaining is to determine the conditions under which “wineglass” wormholes (or quasi-oscillatory wormholes of the type that we describe in section 4) can be the dominant connected saddles of the gravitational path integral, rendering inflation as the most probable outcome for the early Universe cosmology.
@@ 3 The Einstein-scalar-radiation model
Our starting point is the Einstein–Hilbert action with a minimally coupled scalar field $\phi$ (the inflaton), in the presence of gauge/radiation fields which play an important role in shaping the solutions of interest. This is a quite general model that forms a subsector of the Standard Model in the case that the scalar inflaton is identified with the Higgs boson Bezrukov and Shaposhnikov (2008); Barvinsky et al. (2008); De Simone et al. (2009); Barbon and Espinosa (2009); Barvinsky et al. (2009, 2012); Bezrukov et al. (2011, 2015); Betzios et al. (2025); Gialamas et al. (2025). We shall work in four dimensions motivated from phenomenological purposes, but similar solutions can be found in an arbitrary number of dimensions.
@@
where $\mathcal{R}$ denotes the scalar curvature, $V(\phi)$ is the scalar field potential, $\kappa=1/M_{\rm Pl}^{2}$ and $\mathcal{S}_{\rm GHY}$ is the Gibbons-Hawking-York boundary term which has been added to render the variational problem well defined. Higher–order curvature corrections and possible non-minimal couplings between gravity and the scalar field are also expected to be generated at the quantum level. Such terms could, in principle, have a significant impact on the inflationary dynamics of the system, for example by flattening the scalar potential and modifying the inflationary trajectory leading to different inflationary parameters (see Bezrukov et al. (2011, 2015); Gialamas et al. (2025) and Betzios et al. (2025) for an application in the present context). Nevertheless, in what follows we shall keep the discussion and analysis general within our minimal setup, since it can be applied to any particular inflationary model.
@@
The dominance of $SO(4)$-symmetric Euclidean configurations in the path integral Coleman (1985), which reflects the universe’s observed isotropy and homogeneity, motivates the use of the closed (spherical) case, $k=1$ , which we shall adopt in the rest of this work. An important aspect to consider is the sign of the electromagnetic energy density $\tilde{\rho}^{E}_{\rm rad}$. To obtain the Euclidean wormhole solutions for the cosmological purposes of interest Betzios and Papadoulaki (2024a), it is necessary that $\tilde{\rho}^{E}_{\rm rad}<0\,$. This requirement becomes clear from the third equation in (9), where the curvature term $1/a^{2}$ tends to reduce the scale factor, whereas the term $\tilde{\rho}^{E}_{\rm rad}/a^{4}$, with $\tilde{\rho}^{E}_{\rm rad}<0$ , contributes to its expansion. The competition between these terms gives rise to the wormhole neck and supports the subsequent accelerated expansion of the universe. The condition $\tilde{\rho}^{E}_{\rm rad}<0\,$ can be naturally realized if the magnetic field contribution dominates over the electric one in the early universe Betzios et al. (2025) [^9] A similar mechanism can also work in braneworld models such as the ones studied in Antonini and Bariuan (2024) and it would be interesting to extend our analysis in this direction.. In particular, the Euclidean Lagrangian of the electromagnetic field in the $U(1)$ case is $\mathcal{L}_{\rm rad}^{E}=F_{\mu\nu}^{E}F^{E\,\mu\nu}/4=(E^{2}+B^{2})/2\,$, which, through the corresponding electromagnetic energy-momentum tensor, ${T^{E}_{\rm rad}}^{\mu}_{\,\,\nu}$ , leads to a Euclidean radiation energy density of the form
@@
A similar mechanism can be demonstrated for axions (or two form) fields instead of gauge (or one form) fields, see for example Betzios and Papadoulaki (2024a). This again stems from the fact that the axionic contribution to the energy density $\rho_{\rm axion}$ arising from a three form flux $H_{ijk}$ is negative in the Euclidean signature, since it is always of the magnetic type (no $\tau$ component). A final type of mechanism giving rise to Euclidean wormhole saddles is to consider slices that have negative curvature i.e. k=1k=-1\, (combined now with dominant electric radiation density). The slices should be then compactified (in analogy to higher genus Riemann surfaces in two dimensions), if the wormhole interpretation of the backgrounds is to be retained Betzios et al. (2019).